# Meta4XNLI: A Crosslingual Parallel Corpus for Metaphor Detection and Interpretation

### v1

@@ Abstract
Metaphors, although occasionally unperceived, are ubiquitous in our everyday language. Thus, it is crucial for Language Models to be able to grasp the underlying meaning of this kind of figurative language. In this work, we present Meta4XNLI, a novel parallel dataset for the tasks of metaphor detection and interpretation that contains metaphor annotations in both Spanish and English. We investigate language models’ metaphor identification and understanding abilities through a series of monolingual and cross-lingual experiments by leveraging our proposed corpus. In order to comprehend how these non-literal expressions affect models’ performance, we look over the results and perform an error analysis. Additionally, parallel data offers many potential opportunities to investigate metaphor transferability between these languages and the impact of translation on the development of multilingual annotated resources.
@@ 1 Introduction
Metaphor is commonly characterized as the understanding of an abstract concept in terms of another from a more concrete domain. According to Lakoff and Johnson (1980), we can establish a distinction between conceptual metaphors, cognitive mappings that arise from the association between source and target domains, and linguistic metaphors, the expression of these mappings through language. The pervasiveness of metaphors in our daily speech makes it fundamental for language models to be able to process them accordingly, in order to achieve a satisfactory interaction between users and these tools. In addition, metaphor processing may have implications for other Natural Language Processing (NLP) tasks such as Machine Translation Mao, Lin, and Guerin (2018); Schäffner (2004); Shutova, Teufel, and Korhonen (2013), political discourse analysis Charteris-Black (2011); Prabhakaran, Rei, and Shutova (2021); Rodríguez et al. (2023) or hate speech Lemmens, Markov, and Daelemans (2021), among others. Since in this work we study metaphor occurrence in natural language sentences, we will focus on linguistic metaphors only.
@@
The most explored task so far is metaphor detection or identification, approached as a sequence labeling task grounded on different theoretical proposals Wilks (1975, 1978); Searle (1979); Black (1962). The methodology of most widespread use currently are the MIPVU guidelines Steen et al. (2010), which rely on the mismatch between the basic and contextual meaning of a potential metaphor. The application of this procedure resulted in the publication of the referential dataset VUAM. As usual, most published work is English-centered, although multilingual and cross-lingual approaches are increasingly attaining more popularity. Yet available resources for other languages are still scarce, of reduced size or automatically labeled and non-parallel.
Although metaphor interpretation has been less explored than detection, in the recent years it has arisen a growing interest that is reflected in the celebration of the FigLang 2022 Shared Task on Understanding Figurative Language Saakyan et al. (2022). Previous work commonly approached it as a paraphrasing task Shutova (2010); Shutova, Cruys, and Korhonen (2012); Shutova, Teufel, and Korhonen (2013); Shutova (2013); Bizzoni and Lappin (2018). Most recent works frame it within the task of Natural Language Inference (NLI), which consists in determining the relationship between a premise and a hypothesis, generally entailment, neutral or contradiction. In this way, a series of new datasets were published to explore this line of research.
Most of them are limited to one language, typically English, or consist of premises and hypotheses sentences that are almost identical except for the metaphorical expressions, replaced by their literal or their antonym to construct entailment and contradiction pairs Agerri (2008); Mohler, Tomlinson, and Bracewell (2013); Chakrabarty et al. (2021); Stowe, Utama, and Gurevych (2022); Chakrabarty et al. (2022); Kabra et al. (2023). This mechanism to develop datasets might appear fruitful. Nonetheless, these artifacts do not constitute uses of metaphorical expressions in natural language and might yield to inconclusive results when evaluating the performance of the models. Following the state of the art and to compensate the mentioned shortcomings, in this work we will explore metaphor detection as a sequence labeling task and metaphor interpretation via NLI from a cross-lingual perspective with the aid of our proposed cross-lingual and parallel dataset with metaphor annotations for both tasks. Therefore, our contributions can be listed as follows:
We provide Meta4XNLI (M4X), a collection of existing NLI datasets, XNLI Conneau et al. (2018) and esXNLI Artetxe, Labaka, and Agirre (2020), enriched with metaphor annotations for the tasks of detection and interpretation in Spanish (ES) and English (EN). To the best of our knowledge, this is the first multilingual parallel dataset with metaphor annotations in natural language sentences. Meta4XNLI’s characteristics include:
Parallel data and metaphorical annotations at token and premise-hypothesis pair levels.
Enables cross-lingual analysis of metaphor.
Metaphor annotations on texts of multiple domains and natural language sentences.
Contains texts translated in both directions: EN -> ES, ES -> EN, which enables analysis of translation impact on metaphor.
Monolingual and cross-lingual experiments in various setups by leveraging Meta4XNLI. For metaphor detection, we evaluated with Meta4XNLI models trained on other corpora; we fine-tuned and evaluated Masked Language Models (MLM, encoders only) with Meta4XNLI for both languages and also in zero-shot scenarios. Through the usage of different datasets and a cross-lingual approach, our aim is to explore the generalization capabilities of MLMs and the extent of knowledge transference when it comes to metaphor processing. With respect to metaphor interpretation, we framed the task within NLI. In order to study whether MLMs struggle with metaphorical expressions understanding, on one hand, we tested MLMs fine-tuned for the task of NLI with pairs containing and lacking metaphors from Meta4XNLI. On the other, we trained and evaluated the models on pairs without metaphor instances and the whole dataset.
Code and data publicly available in https://github.com/elisanchez-beep/meta4xnli
@@
[…]
2 Related Work
In this section, we will give an overview of most significant works that focused on metaphor processing. In the first place, we analyse metaphor detection and interpretation developed in EN, since most publications and datasets are English-centered. Afterwards, we will explore multi- and cross-lingual approaches for both tasks.
Metaphor Detection
Initially, most work for the task of metaphor detection was corpus-based Charteris-Black (2004); Semino (2017). However, the growing interest throughout the last years led to the celebration of FigLang shared tasks Leong, Beigman Klebanov, and Shutova (2018); Leong et al. (2020), which promoted multiple approaches to address it as a sequence labeling task via deep learning techniques. The most used methodology consisted in training a model with specific features, either linguistics-related Stowe and Palmer (2018) or of other nature such as abstractness, visual or emotion-related information Tsvetkov et al. (2014); Bizzoni and Ghanimifard (2018); Tong, Shutova, and Lewis (2021); Neidlein, Wiesenbach, and Markert (2020).
The arrival of Transformer-based models Devlin et al. (2019) led to an increase in the overall performance of the task. Most fine-tuned models are founded on linguistic theories, such as MIP (Steen et al., 2010) or Selectional Preference (SP) Wilks (1975); Percy (1958), which, generally speaking, address metaphor as a contrast between basic and contextual meaning. The MIP approach was extended to MIPVU to develop the referential corpus in EN: the VUAM dataset Steen et al. (2010), which covers texts of multiple domains and annotations at token level, and was used for the Shared Tasks along with the release of TOEFL dataset Leong et al. (2020). Other well-known datasets are TroFi Birke and Sarkar (2006), of considerable smaller size and restrained to verbs; or MOH-X dataset Mohammad, Shutova, and Turney (2016), also focusing on metaphorical and literal examples of verbs. Other available corpora covers texts from a single domain, such as tweets Zayed, McCrae, and Buitelaar (2019) or news headlines, in NewsMet dataset Joseph et al. (2023).
The combination of pre-trained language models and these available resources brought forth multiple models fine-tuned for metaphor detection with ad hoc architectures. For instance, Song et al. (2021) propose Mr-BERT, a model able to extract the grammatical and semantic relations of a metaphorical verb and its context. RoPPT Wang et al. (2023) takes into account dependency trees information to extract the terms most relevant to the target word. The purpose of other published models is to identify the metaphoric span of the sentence, namely MelBERT Choi et al. (2021), based on MIP and SP theories, as well as BasicBERT Li et al. (2023a). To alleviate scarcity of metaphor annotated data, CATE Lin et al. (2021) is a ContrAstive Pre-Trained ModEl that uses semi-supervised learning and self-training.
Others use additional linguistic resources besides datasets, like Babieno et al. (2022), that take advantage of Wiktionary definitions to build their MLM MIss RoBERTa WiLDe; FrameBERT Li et al. (2023b) makes use of FrameNet Fillmore, Baker, and Sato (2002) to extract the concept of the detected metaphor. MisNET Zhang and Liu (2022) exploits dictionary resources and is based on linguistic theories to predict metaphors at word level. The model of Wan et al. (2021) learns from glosses of the definition of the contextual meaning of metaphors. Maudslay and Teufel (2022) present what they call a Metaphorical Polysemy Detection model by exploiting WordNet and Word Sense Disambiguation (WSD) to perform the detection. Another approach is to frame metaphor detection within another NLP task: such as Zhang and Liu (2023), who adopt a multi-task learning framework where knowledge from WSD is leveraged to identify metaphors; Feng and Ma (2022) develop An Auto-Augmented Sturcture-aware generative model that approaches metaphor detection as a keywords-extraction task; and the work of Dankin, Bar, and Dershowitz (2022), that explores few-shot scenarios from a Yes-No Question-Answering perspective. Badathala et al. (2023) also propose a multi-task approach to detect metaphor and hyperbole altogether although at sentence level.
Among these examples, the state of the art on the task approached as sequence labelling and evaluated on VUA-20, our experimental setup, is the following: DeBERTa-large (73.79 F1) fine-tuned and evaluated on VUAM Sanchez-Bayona and Agerri (2022), BasicBERT (73.3 F1), FrameBERT (73.0), RoPPT (72.8 F1) and MelBERT (72.3 F1). These scores show the complexity of the task and how there is still room for improvement to achieve competitive performance.
Metaphor Interpretation
The evaluation of metaphor interpretation remains a difficult issue, thus most works frame it within other NLU tasks, namely paraphrasing or NLI. Among the paraphrase current, we can find both supervised Shutova, Teufel, and Korhonen (2013); Shutova (2013) and unsupervised Shutova (2010); Shutova, Cruys, and Korhonen (2012) approaches. The work of Bollegala and Shutova (2013) explores the generation of literal paraphrases for metaphorical verbs unsupervisedly. Bizzoni and Lappin (2018) developed a corpus with sentences that contain metaphorical expressions and a set of literal paraphrases ranked according to their acceptability. They exploit their resource to test deep learning systems that approach metaphor interpretation as a classification and ranking task. Mao, Lin, and Guerin (2021) focused on the paraphrase of verb metaphors as well. They take advantage of MOH-X Mohammad, Shutova, and Turney (2016) and VUAM to test BERT’s capability to generate the most probable literal substitute. Pedinotti et al. (2021) develop an evaluation dataset in EN with 300 instances that include conventional and novel metaphors, as well as literal and nonsense sentences. They exploit it to test BERT’s ability to interpret metaphors and discriminate among the different types of sentences, in addition to examine how MLMs encode this knowledge.
[…]
@@ 2 Related Work
Rakshit and Flanigan (2023) introduce FigurativeQA, which gathers 1000 yes/no questions in EN from the reviews domain that include metaphorical and literal context examples in addition to other figurative language phenomena, such as simile, hyperbole, idiom and sarcasm to probe models. Other works like Wachowiak and Gromann (2023); Pitarch, Bernad, and Gracia (2023) center their research on examining whether generative models understand conceptual metaphors and their reasoning capabilities, the former makes use of existing datasets in EN and ES and the latter proposes a new resource only in EN.
@@
A popular approach that continues being explored is the study of metaphor interpretation framed in the NLI or Recognizing Textual Entailment (RTE) task Agerri (2008); Mohler, Tomlinson, and Bracewell (2013). For example, Chakrabarty et al. (2021) propose 12500 instances in EN collected from already existing datasets for RTE, which cover simile, metaphor and irony. Out of the total dataset, only 600 pairs are metaphor related. In addition, their data presents some bias that does not truthfully represent the use of metaphors in natural language. For instance, entailment pairs hypotheses are generated by literal substitution of the metaphorical expression from the premise. This also occurs with non-entailment. In this case, the metaphor from the premise is replaced by an antonym to generate the hypothesis.
@@
The work of Zayed, McCrae, and Buitelaar (2020) aimed at creating a gold standard for metaphor interpretation. Therefore, they developed a dataset of 2500 tweets with definitions of verb-noun metaphorical expressions with the aid of lexical resources and word embeddings. However, they do not present any experiments on models to evaluate their dataset. Stowe, Utama, and Gurevych (2022) introduce IMPLI, a semi-automatic constructed dataset for EN to evaluate RoBERTa-like models performance on figurative language, specifically idioms and metaphors, within the NLI task. Their resource is considerably larger in size (24,029 silver and 1,831 gold sentence pairs) and covers entailment and non-entailment (a merge of neutral and contradiction) relations. Similarly to the previous work, some biases are introduced throughout the generation of this dataset. Entailed pairs only consist of sentences that include a metaphorical expression and the entailed hypothesis is simply its literal paraphrase. On the other hand, non-entailment pairs were developed by “inventing” a literal but not true meaning of the figurative expression. Chakrabarty et al. (2022) propose FLUTE, an explanation-based dataset in EN of 9000 NLI pairs that include sarcasm, simile, metaphor and idioms that they later exploit to evaluate T5 Raffel et al. (2023). FLUTE differs from the other resources in that each entailed/contradicted pair is accompanied by an explanation. However, premises and hypotheses are mainly based on lexical substitution: e.g. His dark clothes had a negative effect in the shadows -> His dark clothes were a plus in the shadows.
@@
All the work we just mentioned is centered in English only, thus most relevant resources developed for metaphor processing comprehend data in this language, both for detection and interpretation tasks. In the case of metaphor detection, in order to compensate the lack of language diversity, some monolingual datasets were published in other languages, although they are commonly of reduced size and still are monolingual, e.g. KOMET corpus Antloga (2020) in Slovene, CoMeta Sanchez-Bayona and Agerri (2022) in Spanish, Estonian Aedmaa, Köper, and Schulte im Walde (2018), German Köper and Schulte im Walde (2016) or Polish adjective-noun metaphors Mykowiecka, Marciniak, and Wawer (2018).
@@
Among the datasets that cover more than one language, we can find LCC Mohammad, Shutova, and Turney (2016): it comprises texts of English, Spanish, Russian and Farsi and provides source/target, metaphoricity degree and metaphor/literal annotations on sentences. However, there are not labels at token level, annotations were extracted automatically, only a subset was manually validated and it is not parallel corpora. The work of Levin et al. (2014) proposes the CCM corpora, which also includes sentences in the same four languages as the LCC dataset. Nevertheless, it is composed of source and target mappings, thus it is centered on conceptual metaphors. Schuster and Markert (2023) explore static embeddings to generate a cross-lingual dataset out of existing resources in German, English and Polish but focused on adjective-noun metaphor pairs and non-parallel texts. Similarly, Berger (2022) explores transfer learning techniques, such as neural machine translation, cross-lingual word embeddings or multilingual pre-trained models to obtain a dataset for metaphor detection in German from English corpora.
@@
In addition to data generation, experimental settings of many recent publications shifted to multi- and/or cross-lingual approaches, which aim at analysing metaphor transferability among different languages. Previous to pre-trained language models, Tsvetkov et al. (2014) already explored unsupervised methods to evaluate metaphor identification of specific syntactic constructions in English, Spanish, Russian and Farsi. Shutova et al. (2017) as well experimented with semi- and unsupervised learning, specifically with clustering techniques for metaphor detection in English, Spanish and Russian. Aghazadeh, Fayyaz, and Yaghoobzadeh (2022) investigate whether pre-trained language models are able to encode metaphorical meaning through the task of metaphor detection in English, Farsi, Russian and Spanish. Not only the trend is to mix languages but also types of figurative language. The work of Lai, Toral, and Nissim (2023) presents a combination of joint-models that are able to detect if a sentence has hyperbole, idioms or metaphors in these same four languages, nevertheless, it does not identify the figurative language span.
@@
With respect to metaphor interpretation, the range of available datasets in multiple languages is more limited than that for detection. Most corpora just mentioned only concerns EN: FigurativeQA Rakshit and Flanigan (2023), Fig-QA Liu et al. (2022), IMPLI Stowe, Utama, and Gurevych (2022) or FLUTE Chakrabarty et al. (2022). The exception is the work of Kabra et al. (2023), who delve into figurative language understanding from a multi-lingual and multi-cultural perspective. They present MABL, a dataset for seven languages with a high number of speakers, such as Hindi, Swahili, or Sudanese. They demonstrate that socio-cultural features pose an essential impact on the conceptual mappings that are later materialized in linguistic metaphors. They exploit it to evaluate language models and provide an insight of how English- or Western-centered the training process of these models can be.
This overview gives an account of the prevalence of EN in the research of metaphor processing. Monolingual datasets for detection in other languages are of a very reduced size in comparison to those in EN. Other limitations of metaphor detection corpora is that datasets with a larger amount of samples do not include annotations at token level or are limited to metaphorical expressions of specific POS. Namely LCC dataset, which is the most utilised multilingual dataset for the task, however, its original annotations are developed at sentence level and it does not contain parallel text.

### v2

# Meta4XNLI: A Cross-lingual Parallel Corpus for Metaphor Detection and Interpretation

@@ Abstract
Metaphors are a ubiquitous but often overlooked part of everyday language. As a complex cognitive-linguistic phenomenon, they provide a valuable means to evaluate whether language models can capture deeper aspects of meaning, including semantic, pragmatic, and cultural context. In this work, we present Meta4XNLI, the first parallel dataset for Natural Language Inference (NLI) newly annotated for metaphor detection and interpretation in both English and Spanish. Meta4XNLI facilitates the comparison of encoder- and decoder-based models in detecting and understanding metaphorical language in multilingual and cross-lingual settings. Our results show that fine-tuned encoders outperform decoders-only LLMs in metaphor detection. Metaphor interpretation is evaluated via the NLI framework with comparable performance of masked and autoregressive models, which notably decreases when the inference is affected by metaphorical language. Our study also finds that translation plays an important role in the preservation or loss of metaphors across languages, introducing shifts that might impact metaphor occurrence and model performance. These findings underscore the importance of resources like Meta4XNLI for advancing the analysis of the capabilities of language models and improving our understanding of metaphor processing across languages. Furthermore, the dataset offers previously unavailable opportunities to investigate metaphor interpretation, cross-lingual metaphor transferability, and the impact of translation on the development of multilingual annotated resources.
@@ 1 Introduction
Metaphor is commonly characterized as the understanding of an abstract concept in terms of another concept from a more concrete domain. According to Lakoff and Johnson (1980), we can establish a distinction between conceptual metaphors, cognitive mappings that arise from the association between source and target domains, and linguistic metaphors, the expression of these mappings through language. The pervasiveness of metaphors in our daily speech makes it fundamental for language models to be able to process them accordingly, in order to achieve a satisfactory interaction between users and these tools. In addition, metaphor processing may have implications for other Natural Language Processing (NLP) tasks such as Machine Translation Mao, Lin, and Guerin (2018); Schäffner (2004); Shutova, Teufel, and Korhonen (2013), political discourse analysis Charteris-Black (2011); Prabhakaran, Rei, and Shutova (2021); Rodríguez et al. (2023) or hate speech Lemmens, Markov, and Daelemans (2021), among others. Since we study metaphor occurrence in natural language sentences, in this work, we will focus on linguistic metaphors only.
@@
The task most explored to date is metaphor detection or identification, typically framed as a sequence labeling problem grounded on different theoretical proposals Wilks (1975, 1978); Searle (1979); Black (1962). The most popular methodology is perhaps the one defined by the MIPVU guidelines Steen et al. (2010), which relies on the mismatch between the basic and contextual meaning of a potential metaphor. The application of this procedure resulted in the publication of the referential dataset VUAM. As usual, most published work is English-centered, although multilingual and cross-lingual approaches are increasingly gaining popularity. However, resources for other languages are still scarce, of reduced size, or automatically labeled and non-parallel.
@@
| H3: She knew exactly what we were talking about. | Contra | No |
Although metaphor interpretation has been less explored than detection, there is a growing interest in recent years, reflected, for example, in the celebration of the FigLang 2022 Shared Task on Understanding Figurative Language Saakyan et al. (2022). Previous work commonly approached understanding as a paraphrasing task Shutova (2010); Shutova, Cruys, and Korhonen (2012); Shutova, Teufel, and Korhonen (2013); Shutova (2013); Bizzoni and Lappin (2018). However, most recent works frame it within the task of Natural Language Inference (NLI), which consists of determining the relationship between a premise and a hypothesis, generally entailment, neutral, or contradiction Agerri (2008); Chakrabarty et al. (2021); Stowe, Utama, and Gurevych (2022); Chakrabarty et al. (2022); Kabra et al. (2023).
Nevertheless, most of the datasets published in this particular paradigm are limited to one language, typically English, or consist of premises and hypotheses that are identical except for the metaphorical expressions which are replaced by their literal or antonym counterparts to construct entailment and contradiction pairs Agerri (2008); Mohler, Tomlinson, and Bracewell (2013); Chakrabarty et al. (2021); Stowe, Utama, and Gurevych (2022); Chakrabarty et al. (2022); Kabra et al. (2023). Although this lexical substitution mechanism to develop datasets might appear fruitful, these artifacts do not represent uses of metaphorical expressions in natural language occurring text.
This work addresses these shortcomings by exploring metaphor detection as a sequence labeling task and metaphor interpretation via NLI from a multilingual and cross-lingual perspective by leveraging a parallel dataset with newly added annotations for both tasks. Our contributions can be summarized as follows:
We provide Meta4XNLI (M4X), a collection of existing NLI datasets, XNLI Conneau et al. (2018) and esXNLI Artetxe, Labaka, and Agirre (2020), enriched with metaphor annotations for the tasks of detection and interpretation in Spanish (ES) and English (EN). To the best of our knowledge, this is the first multilingual parallel dataset with metaphor annotations for both tasks in naturally occurring text. Meta4XNLI’s main features include:
13K parallel sentences with metaphorical annotations at the token and premise-hypothesis pair levels. An example of our dataset is shown in Figure 1.
Supports cross-lingual analysis of metaphorical language use and transfer.
Metaphor annotations across texts of multiple domains and natural language sentences.
Enables the study of how metaphorical expressions influence a natural language understanding task, such as NLI, providing a foundation for evaluating the metaphor interpretation abilities of language models.
Contains texts translated in both directions: EN -> ES, ES -> EN, facilitating analysis of how translation affects metaphor preservation and interpretation.
Monolingual and cross-lingual experiments in various evaluation setups by leveraging Meta4XNLI:
For metaphor detection, we trained models in other datasets and evaluated them cross-domain in Meta4XNLI; we fine-tuned and evaluated Masked Language Models (MLM, encoders only) and decoders-only Large Language Models (LLMs) with Meta4XNLI for both languages and also in zero-shot scenarios. Through the usage of different datasets and a cross-lingual approach, we aim to explore the generalization capabilities of Language Models (LM) and the extent of knowledge transfer when it comes to metaphor processing. Our results show a more competitive performance of MLMs, while LLMs struggle with the task.
Concerning metaphor interpretation, we framed the task within NLI to study the capabilities of Language Models (LMs to refer to both MLMs and LLMs) to understand metaphorical language. First, we tested MLMs and LLMs fine-tuned for the task of NLI with metaphorical and non-metaphorical pairs from Meta4XNLI. Second, we also trained and evaluated MLMs on the whole dataset. In contrast to previous work Rakshit and Flanigan (2023); Stowe, Utama, and Gurevych (2022); Chakrabarty et al. (2022), our results demonstrate that LMs obtain worse results in NLI whenever metaphorical expressions affect the inferential task.
Code and data publicly available111https://huggingface.co/datasets/HiTZ/meta4xnli.
2 Related Work
In this section, we present an overview of the most significant works focused on metaphor processing. First, we analyze metaphor detection and interpretation developed in EN, since most previous works are English-centered. Afterwards, we discuss multi- and cross-lingual approaches for both tasks.
Metaphor Detection
Initially, the majority of the work on metaphor detection was corpus-based Charteris-Black (2004); Semino (2017). However, the growing interest over the last years led to the celebration of FigLang shared tasks Leong, Chee Wee and Beigman Klebanov, Beata and Shutova, Ekaterina(2018) (Ben); Leong et al. (2020), which promoted multiple approaches to address it as a sequence labeling task addressed with deep learning techniques. The most popular methodology consisted of training a model with specific features, either linguistically related Stowe and Palmer (2018) or of other nature, such as abstractness, visual, or emotion-related information Tsvetkov et al. (2014); Bizzoni and Ghanimifard (2018); Tong, Shutova, and Lewis (2021); Neidlein, Wiesenbach, and Markert (2020).
The arrival of Transformer-based models Devlin et al. (2019) led to huge improvements in this task. Most fine-tuned models are founded on linguistic theories, such as MIP (Steen et al., 2010) or Selectional Preference (SP) Wilks (1975); Percy (1958), which, generally speaking, address metaphor as a contrast between basic and contextual meaning. The MIP approach was extended to MIPVU to develop the reference corpus in EN: the VUAM dataset Steen et al. (2010), which covers texts of multiple domains and annotations at token level, and was used for the shared tasks along with the release of TOEFL dataset Leong et al. (2020). Other well-known datasets are TroFi Birke and Sarkar (2006), considerably smaller in size and restricted to verbs; or the MOH-X dataset Mohammad, Shutova, and Turney (2016), which also focuses on metaphorical and literal examples of verbs. Other available corpora cover texts from a single domain, such as tweets Zayed, McCrae, and Buitelaar (2019) or news headlines, in NewsMet dataset Joseph et al. (2023).
The combination of pre-trained language models and these available resources brought forth multiple models fine-tuned for metaphor detection with ad hoc architectures. For instance, Song et al. (2021) propose Mr-BERT, a model capable of extracting the grammatical and semantic relations of a metaphorical verb and its context. RoPPT Wang et al. (2023) takes into account information from dependency trees to extract the terms most relevant to the target word. The purpose of other published models is to identify the metaphoric span of the sentence, namely MelBERT Choi et al. (2021), based on MIP and SP theories, as well as BasicBERT Li et al. (2023a). To alleviate the scarcity of metaphor-annotated data, CATE Lin et al. (2021) is a ContrAstive Pre-Trained ModEl that uses semi-supervised learning and self-training.
Others use additional linguistic resources besides datasets, like Babieno et al. (2022), which takes advantage of Wiktionary definitions to build their MLM MIss RoBERTa WiLDe; FrameBERT Li et al. (2023b) uses FrameNet Fillmore, Baker, and Sato (2002) to extract the concept of the detected metaphor. MisNET Zhang and Liu (2022) exploits dictionary resources and is based on linguistic theories to predict word-level metaphors. The model of Wan et al. (2021) learns from glosses of the definition of the contextual meaning of metaphors. Maudslay and Teufel (2022) present what they call a Metaphorical Polysemy Detection model by exploiting WordNet and Word Sense Disambiguation (WSD) to perform the detection. Another approach is to frame metaphor detection within another NLP task, such as Zhang and Liu (2023), who adopt a multi-task learning framework where knowledge from WSD is leveraged to identify metaphors; Feng and Ma (2022) apply an Auto-Augmented Structure-aware generative model that approaches metaphor detection as a keywords-extraction task; and the work of Dankin, Bar, and Dershowitz (2022), which explores few-shot scenarios from a Yes-No Question-Answering perspective. Badathala et al. (2023) also propose a multi-task approach to detect metaphor and hyperbole, although at the sentence level.
Among these examples, the state of the art on the task evaluated as sequence labeling on VUA-20 was the following: DeBERTa-large (73.79 F1) fine-tuned and evaluated on VUAM Sanchez-Bayona and Agerri (2022), BasicBERT (73.3 F1), FrameBERT (73.0), RoPPT (72.8 F1) and MelBERT (72.3 F1). These scores show the complexity of the task and how there is still room for improvement to achieve competitive performance.
Metaphor Interpretation
The evaluation of metaphor interpretation remains a difficult, open research problem, which is why most works frame it within other NLU tasks, namely, paraphrasing or NLI. Among the works based on paraphrasing, we can find both supervised Shutova, Teufel, and Korhonen (2013); Shutova (2013) and unsupervised Shutova (2010); Shutova, Cruys, and Korhonen (2012) approaches. The work of Bollegala and Shutova (2013) explores the generation of literal paraphrases for metaphorical verbs in an unsupervised manner. Bizzoni and Lappin (2018) developed a corpus with sentences that contain metaphorical expressions and a set of literal paraphrases ranked according to their acceptability. They exploit their resource to test deep learning systems that approach metaphor interpretation as a classification and ranking task. Mao, Lin, and Guerin (2021) also focused on paraphrasing of verbal metaphors as well. They take advantage of MOH-X Mohammad, Shutova, and Turney (2016) and VUAM to test BERT’s capability to generate the most probable literal substitute. Pedinotti et al. (2021) provide an evaluation dataset in EN with 300 instances that include conventional and novel metaphors, as well as literal and nonsense sentences. They exploit it to test BERT’s ability to interpret metaphors and discriminate among the different types of sentences, in addition to examining how MLMs encode this knowledge.
@@ 2 Related Work
Rakshit and Flanigan (2023) introduce FigurativeQA, which gathers 1000 yes/no questions in EN from the reviews domain that include metaphorical and literal examples in addition to other figurative language phenomena, such as simile, hyperbole, idiom, and sarcasm to probe models. Other works like Wachowiak and Gromann (2023); Pitarch, Bernad, and Gracia (2023) center their research on examining whether generative models understand conceptual metaphors and their reasoning capabilities.
@@
A popular approach that continues to be explored is the study of metaphor interpretation framed in the NLI or Recognizing Textual Entailment (RTE) task Agerri (2008); Mohler, Tomlinson, and Bracewell (2013). For example, Chakrabarty et al. (2021) propose 12500 instances in EN collected from existing datasets for RTE, which cover simile, metaphor, and irony. Of the total dataset, only 600 pairs are metaphor-related. In addition, their data is generated by lexical substitution, which does not truthfully represent the use of metaphors in natural language. For instance, entailment pairs of hypotheses are generated by the literal substitution of the metaphorical expression from the premise. This also occurs with non-entailment. In this case, the metaphor from the premise is replaced by an antonym to generate the hypothesis.
@@
The work of Zayed, McCrae, and Buitelaar (2020) aimed to create a gold standard for metaphor interpretation. They developed a dataset of 2500 tweets with definitions of verb-noun metaphorical expressions with the aid of lexical resources and word embeddings. However, they do not present any experiments on models to evaluate their dataset. Stowe, Utama, and Gurevych (2022) introduce IMPLI, a semi-automatic constructed dataset for EN to evaluate RoBERTa-like models’ performance on figurative language, specifically idioms and metaphors, using NLI as an evaluation framework. Their resource is considerably larger (24,029 silver and 1,831 gold sentence pairs) and covers entailment and non-entailment (a merge of neutral and contradiction) relations.
As is the case with other previously reviewed datasets, some biases are introduced during the generation of this dataset. Entailed pairs only consist of sentences that include a metaphorical expression, while the entailed hypothesis corresponds to its literal paraphrase. Furthermore, non-entailment pairs were developed by “creating” a literal but not natural meaning of the figurative expression. Chakrabarty et al. (2022) propose FLUTE, an explanation-based dataset in EN of 9000 NLI pairs that include sarcasm, simile, metaphor, and idioms. FLUTE differs from the other resources in that each entailed/contradicted pair is accompanied by an explanation. However, premises and hypotheses are based on lexical substitution: e.g., His dark clothes had a negative effect in the shadows -> His dark clothes were a plus in the shadows.
The limited availability of metaphor interpretation datasets in languages other than English, combined with the templatic structure of existing resources, presents a significant obstacle to a generalizable evaluation of the capabilities of language models to understand metaphorical language. Many current datasets rely on lexical substitution methods, which can introduce artifacts that might favor model performance Artetxe, Labaka, and Agirre (2020); Naik et al. (2018). To advance metaphor understanding in multilingual contexts and balance the prevalence of English resources, it is necessary to develop parallel datasets with metaphor annotations in naturally occurring text across more than one language.
@@
The related work discussed so far is centered on English only. In the case of metaphor detection, and to compensate for the lack of language diversity, some monolingual datasets have been published in other languages. However, their size is rather limited and they tend to be monolingual, e.g., the KOMET corpus Antloga (2020) in Slovene, CoMeta Sanchez-Bayona and Agerri (2022) in Spanish, Estonian Aedmaa, Köper, and Schulte im Walde (2018), German Köper and Schulte im Walde (2016) or Polish adjective-noun metaphors Mykowiecka, Marciniak, and Wawer (2018).
@@
Among the datasets that cover more than one language, we can find LCC Mohammad, Shutova, and Turney (2016): it comprises texts of English, Spanish, Russian, and Farsi and provides source/target, metaphoricity degree, and metaphor/literal annotations on sentences. However, there are no labels at the token level, annotations were extracted automatically, only a subset was manually validated, and it is not parallel text. The work of Levin et al. (2014) proposes the CCM corpora, which also includes sentences in the same four languages as the LCC dataset. Nevertheless, it is composed of source and target mappings. Thus, it is centered on conceptual metaphors. Schuster and Markert (2023) explores static embeddings to generate a cross-lingual dataset out of existing resources in German, English, and Polish, but focuses on adjective-noun metaphor pairs and non-parallel texts. Similarly, Berger (2022) explores transfer learning techniques, such as neural machine translation, cross-lingual word embeddings, or multilingual pre-trained language models, to obtain a dataset for metaphor detection in German from English corpora. The work of Wang et al. (2024) presents a parallel dataset in Chinese and English. However, it is not annotated for any of the tasks of our interest.
@@
In addition to data generation, the experimental settings of many recent publications have shifted to multi- and/or cross-lingual approaches, with the objective of analysing metaphor transferability among different languages. Previous to the appearance of pre-trained language models, Tsvetkov et al. (2014) already explored unsupervised methods to evaluate metaphor identification of specific syntactic constructions in English, Spanish, Russian, and Farsi. Shutova et al. (2017) also experimented with semi- and unsupervised learning, specifically with clustering techniques for metaphor detection in English, Spanish, and Russian. Aghazadeh, Fayyaz, and Yaghoobzadeh (2022) investigate whether pre-trained language models are able to encode metaphorical meaning through the task of metaphor detection in English, Farsi, Russian, and Spanish.
The work of Lai, Toral, and Nissim (2023) presents a combination of joint models able to detect whether a sentence contains hyperbole, idioms, or metaphorical expressions in these same four languages. Nevertheless, it does not provide annotations for the text at span level.
@@
Concerning metaphor interpretation, the range of available datasets in multiple languages is more limited than for detection. Most of the data already mentioned is available only in English: FigurativeQA Rakshit and Flanigan (2023), Fig-QA Liu et al. (2022), IMPLI Stowe, Utama, and Gurevych (2022), or FLUTE Chakrabarty et al. (2022). The exception is the work of Kabra et al. (2023), who address figurative language understanding from a multi-lingual and multi-cultural perspective. They present MABL, a dataset for seven languages with a high number of speakers, such as Hindi, Swahili, or Sudanese. They demonstrate that socio-cultural features essentially impact conceptual mappings that later materialize in linguistic metaphors. They evaluate language models on MABL and provide some insights regarding the English- or Western-centered training process of these models.
To summarize, this overview gives an account of the prevalence of EN in metaphor processing research. While some monolingual detection datasets exist for other languages, they are quite small when compared to those in EN. In fact, any fairly sized corpora for metaphor detection different from English do not include annotations at the token level or are limited to metaphorical expressions of a given POS. An example is the LCC dataset, which is a popular non-parallel dataset that only includes annotations at the sentence level.